OpenStudy (anonymous):

I'm confused with this. Can someone PLEASE help? When a ball is tossed into the air, the height in feet at time (t) in seconds is -16t^2+64t. At what time (t) will the ball reach it's highest point?

OpenStudy (anonymous):

I really need help someone... :(

OpenStudy (anonymous):

Set -16t^2+64t = 0 and solve for t (move each term over to the zero side) 16t^2 - 64t = 0; now factor it. 16t (t - 4) = 0 when you solve for t : 16t = 0 and t - 4 = 0, you get t = 0 & 4 those are the two times the ball leaves the ground and come back to the ground, at 0 and at 4 seconds. It will be at its highest point halfway between 0 and 4.